OpenStudy (anonymous):

Please help! I have a word problem that I need to get an equation out of in order to solve with elimination, but I have no idea how to. The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1150. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many one-step equations eliminated today? I know how to solve a system of equations by elimination, I just can't figure out what the system of equations is to solve!

OpenStudy (anonymous):

something random

terenzreignz (terenzreignz):

Okay, first thing to do when given a word problem is representing (relevant) values. Judging by your problem, the relevant values are the number of one-step and two-step equations... right?

OpenStudy (anonymous):

I think so

terenzreignz (terenzreignz):

Okay, so why don't we let x be the number of one-step equations and y be the number of two-step equations?

OpenStudy (anonymous):

alright

terenzreignz (terenzreignz):

Copying and pasting a sentence from your problem... "The number of one-step equations and two-step equations who have been eliminated today is equal to 1150." What does this tell you?

OpenStudy (anonymous):

both of the equations equal 1150?

terenzreignz (terenzreignz):

Well, the number of one step equations is x... ........the number of two step equations is y... and there are 1150 total equations... So can you translate that into Maths-Language? :D

OpenStudy (anonymous):

x + y = 1150? Haha I'm sorry if that's wrong..

terenzreignz (terenzreignz):

Yeah... but it isn't. That's actually quite correct :P

OpenStudy (anonymous):

I thought I was looking for a system of equations though>

OpenStudy (anonymous):

?*

terenzreignz (terenzreignz):

So you have one equation x + y = 1150 We need another... allow me to copy and paste another sentence... " If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300..."

terenzreignz (terenzreignz):

That is to say, we need TWO equations to make a system... we're in the process of finding the second equation now...

OpenStudy (anonymous):

3x - 2y = 1300? Just a guess for now

terenzreignz (terenzreignz):

You're guesses are uncannily good :P That's also correct XD x + y = 1150 3x - 2y = 1300 You have a system, work it out, champ :)

OpenStudy (anonymous):

thank you so much! I really appreciate it!

terenzreignz (terenzreignz):

No problem. :)